An acetate buffer was prepared by mixing 10 mL of 0.1 M acetic acid and 100mL of 0.1 M sodium acetate. What is the pH of the buffer solution?
1
Expert's answer
2021-09-06T13:50:02-0400

Millimoles of CH3COOH=0.1×10=1.0


Millimoles of CH3COONa=0.1×100=10


From, Henderson Hasselbalch equation,

pH=ppKa log "\frac{conjugate base}{acid}"

= 4.24 + log "\frac{20}{1}"

4.24 + 1.3

= 4.54